# Iterated INLA for State and Parameter Estimation in Nonlinear Dynamical Systems

### v1

@@ 5 Discussion and Conclusion
In this paper, we proposed a novel data assimilation algorithm based on the INLA methodology that effectively learns the state and the parameters in nonlinear dynamical systems from data that is sparsely distributed. This is achieved by iteratively linearising the dynamical model, resulting in a GMRF, where one can apply INLA to infer the state and parameters. We prove that this is approximately identical to the Gauß-Newton method for minimising the 4D-Var loss, and demonstrate experimentally that it is numerically robust; it also produces accurate non-Gaussian estimates of the latent variables. Issues remain regarding the scalability of the method: INLA is typically employed for moderately-sized problems in two or three physical dimensions. It is difficult to see this being used in very large-scale applications, such as numerical weather forecasting. We also do not consider non-Gaussian likelihoods here, although this should be a straightforward extension by adopting nested Laplace approximations. We also have not exploited the Markovian structure in the temporal component à la filtering/smoothing, which may help to speed up the algorithm. While the results are promising for the toy models considered here, further investigation is necessary to determine how our method fares in realistic medium-scale scenarios such as optical tomography Arridge and Schotland [2009] and nuclear fusion control [Morishita et al., 2024].
@@
Conceptualisation: ST; Methodology: ST, RA; Software: RA; Writing - original draft: ST; Writing - Review and Editing: RA, MD; Supervision: ST, MD. All authors approved the final submitted draft.
@@ A.1 Sparse linear solve and matrix inversion
$\begin{split}&=\sum_{i=1}^{|\boldsymbol{\theta}|}\log p(\theta_{i})+\frac{1}{2}\log|\mathbf{Q}_{\boldsymbol{u}}(\boldsymbol{\theta})|-(\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})-\frac{1}{2}\boldsymbol{\mu}_{\boldsymbol{u}}(\boldsymbol{\theta}))^{\top}\mathbf{Q}_{\boldsymbol{u}}(\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})(\boldsymbol{\theta})-\boldsymbol{\mu}_{\boldsymbol{u}}(\boldsymbol{\theta}))-\frac{M}{2}\log 2\pi\\[1.0pt] &\quad+\frac{1}{2}\log|\mathbf{R}^{-1}|-(\boldsymbol{y}-\mathbf{H}\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}})^{\top}\mathbf{R}^{-1}(\boldsymbol{y}-\mathbf{H}\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}})-N\log 2\pi\\[4.0pt] &\qquad+\frac{1}{2}\log|\mathbf{Q}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})|-M\log 2\pi+\text{const.},\end{split}$ (46)
@@ Appendix B Proofs and discussions of results
The damped update of the linearisation point $\boldsymbol{u}_{0}^{(n)}$ in (30) is equivalent to minimising the weak-constraint 4D-Var cost (10) using Gauß–Newton.
@@
The Gauß–Newton iteration for minimising (52) reads
@@
is the preconditioner, given by the Gauß–Newton approximation to the Hessian of $J$. Next, denoting
@@
we manipulate the above expression for the Gauß–Newton iteration as follows
@@
Updating the linearisation point $\boldsymbol{u}^{(n)}_{0}$ according to equations (36)–(39) is an approximate Gauß–Newton method for minimising the parameter-averaged 4D-Var cost $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}[-\log p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})]$.
@@
The Gauß–Newton iteration for minimising the cost (74) reads
@@
is the preconditioner, given by the Gauß–Newton approximation of the Hessian of $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}\big{[}J_{\boldsymbol{\theta}}[\boldsymbol{u}]\big{]}$. Now by a similar calculation to that in the proof of Proposition 3.2, one can check that
@@ B.3 Further discussion of the results
which we refer to as the Gauß-Newton approximation of the Hessian. The only difference with the true Hessian is the term $\nabla^{2}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]\bar{\mathbf{Q}}^{-1}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$, which is small if $\nabla^{2}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$ or $\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$ is small. The former holds if the dynamics is weakly nonlinear and the latter holds if $\boldsymbol{u}^{(\infty)}_{0}$ is close to the solution of the deterministic system starting from $\boldsymbol{u}_{0}=\boldsymbol{u}_{b}$ (see (56)). Now, the Laplace approximation for the distribution $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$ is given by
@@
Hence, assuming that $\mathbf{P}^{(\infty)}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})\approx\nabla^{2}J[\boldsymbol{u}_{*}]$, we see that the outputs of Algorithm 1 can be interpreted as the marginals of a “Gauß-Newton-Laplace” approximation of the posterior $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$.
@@
The assumption that $\mathcal{KL}(p(\boldsymbol{\theta}|\boldsymbol{y})||p(\boldsymbol{\theta}|\boldsymbol{u}_{*},\boldsymbol{y}))\approx 0$ should hold if for example $p(\boldsymbol{\theta}|\boldsymbol{y})$ is very peaked and depends weakly on $\boldsymbol{u}$. In this case, the MAP estimate of $p(\boldsymbol{\theta}|\boldsymbol{y})$ should be reliably approximated by $\boldsymbol{u}_{0}^{(\infty)}$.
@@ C.2.4 Baseline details
For our SMC method, we split the procedure into two parts. The first part samples the parameters $\boldsymbol{\theta}_{k}\sim p(\boldsymbol{\theta}|\boldsymbol{y})$ using the particle marginal Metropolis-Hastings (PMMH) method in Andrieu et al. [2010]. This is a Metropolis-Hastings algorithm that approximates the term $p(\boldsymbol{y}|\boldsymbol{\theta})$ in the acceptance ratio $p(\boldsymbol{y}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}^{\prime})/p(\boldsymbol{y}|\boldsymbol{\theta}_{n})p(\boldsymbol{\theta}_{n})$ using a bootstrap particle filter with fixed $\boldsymbol{\theta}$. We used $1,000$ particles for the bootstrap filter and sampled $10,000$ parameters from $p(\boldsymbol{\theta}|\boldsymbol{y})$. We used a burn-in period of $1,000$. In the next step, we used the generated parameters $\boldsymbol{\theta}_{k}$ and computed a sample $\boldsymbol{u}_{k}$ of $p(\boldsymbol{u}|\boldsymbol{\theta}_{k},\boldsymbol{y})$ for each $k$ using the bootstrap particle smoother [Chopin et al., 2020], which are precisely the samples of $p(\boldsymbol{u}|\boldsymbol{y})$. The Euler-Maruyama scheme was used to simulate the dynamics (97) with a time step of $\Delta t=0.01$. For the implementation of PMMH and the bootstrap particle filter/smoother, we used the python package particles [^1] https://github.com/nchopin/particles.
@@ Appendix D Discretisation details
All discretisations are performed using finite differences with the python package findiff [Baer, 2018]. We discretised the linearised operators (106), (117), (125) and (130) using second-order central finite differences, treating the spatial and temporal variables on the same footing (i.e., we do not consider forward time-stepping methods). We imposed appropriate boundary conditions depending on the problem. findiff implements Dirichlet and Neumann boundary conditions. However, some of the experiments require periodic boundary conditions, hence we added this functionality in our fork of the package [^4] https://github.com/rafaelanderka/findiff. On the temporal boundaries, findiff by default uses a forward discretisation of the derivative at the initial time and backward discretisation at the final time, if the conditions are not specified. While this does not make sense physically, we found that using this default set up gave us sufficiently good results for our purpose. The random initial conditions were specified by additional likelihoods at the initial time, where we placed pseudo-observations $p(\boldsymbol{y}|\boldsymbol{u}_{0})$ at time $t=0$ to mimic the initial condition prior $p(\boldsymbol{u}_{0})=\mathcal{N}(\boldsymbol{u}_{0}|\boldsymbol{u}_{b},\mathbf{C})$. Mathematically, this is not a problem by taking $\left.p(\boldsymbol{y}|\boldsymbol{u}_{0})\right|_{\boldsymbol{y}=\boldsymbol{u}_{b}}=\left.\mathcal{N}(\boldsymbol{y}|\boldsymbol{u}_{b},\mathbf{C})\right|_{\boldsymbol{y}=\boldsymbol{u}_{b}}$ as it results in the same posterior distribution. In the future, it might be interesting to encode the initial condition prior directly into the model, using proper temporal time-stepping schemes such as the Crank-Nicolson method to discretise parabolic PDEs.

### v2

@@ 5 Discussion and Conclusion
In this paper, we proposed an algorithm based on the INLA methodology that effectively learns the state and the parameters in nonlinear dynamical systems without resorting to expensive MCMC. This is achieved by iteratively linearising the dynamical model, where one can apply INLA to infer the state and parameters. We prove that this is approximately identical to the Gauss-Newton method for minimising the 4D-Var loss, and demonstrate experimentally that it is numerically robust; it also produces accurate non-Gaussian estimates of the latent variables. Issues remain regarding the scalability of the method: INLA is typically employed for moderately-sized problems in two or three physical dimensions. It is difficult to see this being used in very large-scale applications, such as numerical weather forecasting. We also do not consider non-Gaussian likelihoods here, although this should be a straightforward extension by adopting nested Laplace approximations. We also have not exploited the Markovian structure in the temporal component à la filtering/smoothing, which may help to speed up the algorithm. While the results are promising for the toy models considered here, further investigation is necessary to determine how our method fares in realistic medium-scale scenarios such as optical tomography Arridge and Schotland [2009] and nuclear fusion control [Morishita et al., 2024].
{code}
The code accompanying this paper is available at https://github.com/rafaelanderka/iter-inla.
@@
Conceptualisation: ST; Methodology: RA, ST; Software: RA; Writing - original draft: ST; Writing - Review and Editing: RA, MPD; Supervision: MPD, ST. All authors approved the final submitted draft.
@@ A.1 Sparse linear solve and matrix inversion
$\begin{split}&=\sum_{i=1}^{|\boldsymbol{\theta}|}\log p(\theta_{i})+\frac{1}{2}\log|\mathbf{Q}_{\boldsymbol{u}}(\boldsymbol{\theta})|-\frac{1}{2}(\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})-\boldsymbol{\mu}_{\boldsymbol{u}}(\boldsymbol{\theta}))^{\top}\mathbf{Q}_{\boldsymbol{u}}(\boldsymbol{\theta})(\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})-\boldsymbol{\mu}_{\boldsymbol{u}}(\boldsymbol{\theta}))-\frac{M}{2}\log 2\pi\\[1.0pt] &\quad+\frac{1}{2}\log|\mathbf{R}^{-1}|-\frac{1}{2}(\boldsymbol{y}-\mathbf{H}\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta}))^{\top}\mathbf{R}^{-1}(\boldsymbol{y}-\mathbf{H}\boldsymbol{\mu}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta}))-\frac{N}{2}\log 2\pi\\[4.0pt] &\qquad+\frac{1}{2}\log|\mathbf{Q}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})|-\frac{M}{2}\log 2\pi+\text{const.},\end{split}$ (46)
@@ Appendix B Proofs and discussions of results
The damped update of the linearisation point $\boldsymbol{u}_{0}^{(n)}$ in (30) is equivalent to minimising the weak-constraint 4D-Var cost (10) using Gauss–Newton.
@@
The Gauss–Newton iteration for minimising (52) reads
@@
is the preconditioner, given by the Gauss–Newton approximation to the Hessian of $J$. Next, denoting
@@
we manipulate the above expression for the Gauss–Newton iteration as follows
@@
Updating the linearisation point $\boldsymbol{u}^{(n)}_{0}$ according to equations (36)–(39) is an approximate Gauss–Newton method for minimising the parameter-averaged 4D-Var cost $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}[-\log p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})]$.
@@
The Gauss–Newton iteration for minimising the cost (73) reads
@@
is the preconditioner, given by the Gauss–Newton approximation of the Hessian of $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}\big{[}J_{\boldsymbol{\theta}}[\boldsymbol{u}]\big{]}$. Now by a similar calculation to that in the proof of Proposition 3.2, one can check that
@@ B.3 Further discussion of the results
which we refer to as the Gauss-Newton approximation of the Hessian. The only difference with the true Hessian is the term $\nabla^{2}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]\bar{\mathbf{Q}}^{-1}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$, which is small if $\nabla^{2}\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$ or $\mathcal{L}[\boldsymbol{u}^{(\infty)}_{0}]$ is small. The former holds if the dynamics is weakly nonlinear and the latter holds if $\boldsymbol{u}^{(\infty)}_{0}$ is close to the solution of the deterministic system starting from $\boldsymbol{u}_{0}=\boldsymbol{u}_{b}$ (see (55)). Now, the Laplace approximation for the distribution $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$ is given by
@@
Hence, assuming that $\mathbf{P}^{(\infty)}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta})\approx\nabla^{2}J[\boldsymbol{u}_{*}]$, we see that the outputs of Algorithm 1 can be interpreted as the marginals of a “Gauss-Newton-Laplace” approximation of the posterior $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$.
@@
The assumption that $\mathcal{KL}(p(\boldsymbol{\theta}|\boldsymbol{y})||p(\boldsymbol{\theta}|\boldsymbol{u}_{*},\boldsymbol{y}))\approx 0$ should hold if for example $p(\boldsymbol{\theta}|\boldsymbol{y})$ is very peaked and depends weakly on $\boldsymbol{u}$. In this case, the MAP estimate of $p(\boldsymbol{u}|\boldsymbol{y})$ should be reliably approximated by $\boldsymbol{u}_{0}^{(\infty)}$.
@@ C.2.4 Baseline details
For our SMC method, we split the procedure into two parts. The first part samples the parameters $\boldsymbol{\theta}_{k}\sim p(\boldsymbol{\theta}|\boldsymbol{y})$ using the particle marginal Metropolis-Hastings (PMMH) method in Andrieu et al. [2010]. This is a Metropolis-Hastings algorithm that approximates the term $p(\boldsymbol{y}|\boldsymbol{\theta})$ in the acceptance ratio $p(\boldsymbol{y}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}^{\prime})/p(\boldsymbol{y}|\boldsymbol{\theta}_{n})p(\boldsymbol{\theta}_{n})$ using a bootstrap particle filter with fixed $\boldsymbol{\theta}$. We used $1,000$ particles for the bootstrap filter and sampled $10,000$ parameters from $p(\boldsymbol{\theta}|\boldsymbol{y})$. We used a burn-in period of $1,000$. In the next step, we used the generated parameters $\boldsymbol{\theta}_{k}$ and computed a sample $\boldsymbol{u}_{k}$ of $p(\boldsymbol{u}|\boldsymbol{\theta}_{k},\boldsymbol{y})$ for each $k$ using the bootstrap particle smoother [Chopin and Papaspiliopoulos, 2020], which are precisely the samples of $p(\boldsymbol{u}|\boldsymbol{y})$. The Euler-Maruyama scheme was used to simulate the dynamics (96) with a time step of $\Delta t=0.01$. For the implementation of PMMH and the bootstrap particle filter/smoother, we used the python package particles [^1] https://github.com/nchopin/particles.
@@
Iterated Ensemble Kalman Smoother (iEnKS).
For a single iteration of the EnKS, we used the same configuration of EnKS as above. The result is displayed for 10 iterations.
@@ C.3.3 Baseline details
iEnKS.
For a single iteration of the EnKS, we used the same configuration of EnKS as above. We used 30 iterations to produce the final result.
@@ C.4.3 Baseline details
iEnKS.
For a single iteration of the EnKS, we used the same configuration of EnKS as above. We used 30 iterations to produce the final result.
@@ C.5.3 Baseline details
iEnKS.
For a single iteration of the EnKS, we used the same configuration of EnKS as above. We used 30 iterations to produce the final result.
@@ Appendix D Discretisation details
All discretisations are performed using finite differences with the python package findiff [Baer, 2018]. We discretised the linearised operators (105), (116), (124) and (129) using second-order central finite differences, treating the spatial and temporal variables on the same footing (i.e., we do not consider forward time-stepping methods). We imposed appropriate boundary conditions depending on the problem. findiff implements Dirichlet and Neumann boundary conditions. However, some of the experiments require periodic boundary conditions, hence we added this functionality in our fork of the package, which we do not disclose here to preserve anonymity. On the temporal boundaries, findiff by default uses a forward discretisation of the derivative at the initial time and backward discretisation at the final time, if the conditions are not specified. While this does not make sense physically, we found that using this default set up gave us sufficiently good results for our purpose. The random initial conditions were specified by additional likelihoods at the initial time, where we placed pseudo-observations $p(\boldsymbol{y}|\boldsymbol{u}_{0})$ at time $t=0$ to mimic the initial condition prior $p(\boldsymbol{u}_{0})=\mathcal{N}(\boldsymbol{u}_{0}|\boldsymbol{u}_{b},\mathbf{C})$. Mathematically, this is not a problem by taking $\left.p(\boldsymbol{y}|\boldsymbol{u}_{0})\right|_{\boldsymbol{y}=\boldsymbol{u}_{b}}=\left.\mathcal{N}(\boldsymbol{y}|\boldsymbol{u}_{b},\mathbf{C})\right|_{\boldsymbol{y}=\boldsymbol{u}_{b}}$ as it results in the same posterior distribution. In the future, it might be interesting to encode the initial condition prior directly into the model, using proper temporal time-stepping schemes such as the Crank-Nicolson method to discretise parabolic PDEs.